# Six magnetization plateau phases in a kagome antiferromagnet with $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$-type distortion

### v1

@@ III.3 Calculations with other parameters
0.9 $2.70119\pm 0.03941$ $2.94912\pm 0.001358$ 0.20716
1.0 $2.93238\pm 0.02916$ $3\pm 0$ 0.03846
1.1 $3.01029\pm 0.03901$ $3.09162\pm 0.01731$ 0.02501
1.2 $3.13716\pm 0.03533$ $3.20472\pm 0.02336$ 0.00887
1.3 $3.27278\pm 0.03159$ $3.32751\pm 0.02422$0.00108
1.4 $3.41160\pm 0.02822$ $3.45636\pm 0.02326$0.00672
Since the widths of the plateaus and the phase boundaries have been clarified, we next discuss the phase diagram. Figure 7 depicts the phase diagram corresponding to each plateau, identifying a total of six plateau phases. As discussed above, the phase boundaries are at $J_{\rm g}=1.0$. Consequently, there are the six plateau phases, labeled I to VI. For $M/M_{\rm sat}=5/9$ and 7/9, the finite-size scaling is conducted and error bars are included, whereas for $M/M_{\rm sat}=1/3$, results are shown for the $N=36$ system. This is because calculations were only performed for $N=27$ and 36, making it impossible to determine errors in finite-size scaling. In all regions with respect to $J_{\rm g}$, there exists the broad 1/3 magnetization plateau. Additionally, even when including errors from finite-size scaling, the 5/9 and 7/9 plateaus exist for $J_{\rm g}\leq 1.2$. For comparison, the region of the plateau phase in the classical spin system is also depicted in Fig. 7. While the quantum system exhibits six distinct plateau phases, only one plateau phase appears in the classical system. Additionally, the region for the phase II in the quantum system is clearly broader than its counterpart in the classical system. This is due to quantum fluctuations that further stabilize the plateaus.
@@
To investigate the magnetic structures of these six plateau phases, we calculate the spin correlations at $J_{\rm g}=0.9$ and $J_{\rm g}=1.1$. Figure 8 presents the calculation results of the spin correlation functions. For $M/M_{\rm sat}=1/3$, 5/9, and 7/9, the spin structures at $J_{\rm g}=0.9$ and $J_{\rm g}=1.1$ are distinctly different. Examining Figs. 8(a), 8(c), and 8(e), it is clear that at $J_{\rm g}=0.9$, all three plateaus exhibit the MCS. As shown in Figs. 8(b), 8(d), and 8(f), the sites with the lowest value of $\Braket{S_{i}^{z}}$ at $J_{\rm g}=0.9$ change significantly to exhibit the highest value at $J_{\rm g}=1.1$. For $M/M_{\rm{sat}}=1/3$, Fig. 8(b) indicates that, solely focusing on $\Braket{S_{i}^{z}}$, the state is equivalent to the $\sqrt{3}\times\sqrt{3}$ up-up-down state which is classically stable at $J_{\rm g}>1$. However, the $M/M_{\rm{sat}}=5/9$ and 7/9 plateaus do not appear in the classical spin system. Additionally, from the results shown in Fig. 8(d) and Fig. 8(f), it remains unclear why these plateaus stabilize in the quantum spin system. In $J_{1}-J_{2}-J_{3}$ square kagome lattice model similar to the present model, a plateau phase with spontaneous rotational symmetry breaking has been found to exist in addition to the MCS (VBC) at $M/M_{\rm{sat}}=2/3$ [38]. It is possible that the present model stabilizes under similar principles; however, in the current calculations, no breaking of rotational symmetry has been observed. Further calculations with larger system sizes will be necessary, posing a challenge for future research.
@@ IV CONCLUSIONS
The recently discovered spin-1/2 kagome antiferromagnet $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ is predicted to exhibit the Type 1 and Type 2 exchange interaction parameters, as shown in Table 1, according to separate studies. Inspired by this discovery, we investigated the ground state of the spin-1/2 Heisenberg kagome model with three types of exchange interactions in the magnetic field using the Lanczos-type exact diagonalization method. We confirmed the existence of clear $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9 magnetization plateaus under both the Type 1 and Type 2 parameter sets, identified as the MCS. Our calculation results suggest that these plateaus could potentially be verified experimentally through magnetization measurements under the magnetic field of approximately 300 T. Notably, clear 5/9 and 7/9 magnetization plateaus have not yet been observed in experiments, highlighting the importance of additional experimental investigation into $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. Additionally, by performing calculations of the kagome lattice with parameter sets other than the Type 1 and Type 2, we discovered additional plateau phases at $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9. The mechanism behind the 1/3 plateau can be easily explained by analytically solving a three-site system. In contrast, the mechanisms for the 5/9 and 7/9 plateaus could not be clarified in this study. Further experimental and theoretical research is therefore needed to elucidate these phenomena. These findings indicate that further experimental studies on new kagome compounds may successfully observe these new plateaus.

### v3

@@ III.3 Calculations with other parameters
0.9 $2.70\pm 0.04$ $2.95\pm 0.001$ 0.207
1.0 $2.93\pm 0.03$ $3\pm 0$ 0.038
1.1 $3.01\pm 0.04$ $3.09\pm 0.02$ 0.025
1.2 $3.14\pm 0.04$ $3.20\pm 0.02$ 0.009
1.3 $3.27\pm 0.03$ $3.33\pm 0.02$0.001
1.4 $3.41\pm 0.03$ $3.46\pm 0.02$0.007
Since the widths of the plateaus and the phase boundaries have been clarified, we next discuss the phase diagram. Figure 7 depicts the phase diagram corresponding to each plateau, identifying a total of six plateau phases. As discussed above, the phase boundaries are at $J_{\rm g}=1.0$. Consequently, there are the six plateau phases, labeled I to VI. For $M/M_{\rm sat}=5/9$ and 7/9, the finite-size scaling is conducted and error bars are included, whereas for $M/M_{\rm sat}=1/3$, results are shown for the $N=36$ system. This is because calculations were only performed for $N=27$ and 36, making it impossible to determine errors in finite-size scaling. In all regions with respect to $J_{\rm g}$, there exists the broad 1/3 magnetization plateau. Additionally, even when including errors from finite-size scaling, the 5/9 and 7/9 plateaus exist for $J_{\rm g}\leq 1.2$. For comparison, the region of the plateau phase in the classical spin system, calculated using the simulated annealing technique with the Monte Carlo method, is also depicted in Fig. 7. While the quantum system exhibits six distinct plateau phases, only one plateau phase appears in the classical system. Additionally, the region for the phase II in the quantum system is clearly broader than its counterpart in the classical system. This is due to quantum fluctuations that further stabilize the plateaus. The phenomenon of plateau stabilization has been often observed in frustrated systems such as triangular lattices [46, 47, 48, 49, 50].
@@
To investigate the magnetic structures of these six plateau phases, we calculate the spin correlations at $J_{\rm g}=0.9$ and $J_{\rm g}=1.1$. Figure 8 presents the calculation results of the spin correlation functions. For $M/M_{\rm sat}=1/3$, 5/9, and 7/9, the spin structures at $J_{\rm g}=0.9$ and $J_{\rm g}=1.1$ are distinctly different. Examining Figs. 8(a), 8(c), and 8(e), it is clear that at $J_{\rm g}=0.9$, all three plateaus exhibit the MCS. As shown in Figs. 8(b), 8(d), and 8(f), the sites with the lowest value of $\Braket{S_{i}^{z}}$ at $J_{\rm g}=0.9$ change significantly to exhibit the highest value at $J_{\rm g}=1.1$. For $M/M_{\rm{sat}}=1/3$, Fig. 8(b) indicates that, solely focusing on $\Braket{S_{i}^{z}}$, the state is equivalent to the $\sqrt{3}\times\sqrt{3}$ up-up-down state, which is found to be stable for $J_{\rm g}>1$ in our classical calculations. However, the $M/M_{\rm{sat}}=5/9$ and 7/9 plateaus do not appear in the classical spin system. Additionally, from the results shown in Fig. 8(d) and Fig. 8(f), it remains unclear why these plateaus stabilize in the quantum spin system. In $J_{1}-J_{2}-J_{3}$ square kagome lattice model similar to the present model, a plateau phase with spontaneous rotational symmetry breaking has been found to exist in addition to the MCS (VBC) at $M/M_{\rm{sat}}=2/3$ [51]. It is possible that the present model stabilizes under similar principles; however, in the current calculations, no breaking of rotational symmetry has been observed. Further calculations with larger system sizes will be necessary, posing a challenge for future research.
@@ IV CONCLUSIONS
The recently discovered spin-1/2 kagome antiferromagnet $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$ is predicted to exhibit the Type 1 and Type 2 exchange interaction parameters, as shown in Table 1, according to separate studies. Inspired by this discovery, we investigated the ground state of the spin-1/2 Heisenberg kagome model with three types of exchange interactions in the magnetic field using the Lanczos-type exact diagonalization method. We confirmed the existence of clear $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9 magnetization plateaus under both the Type 1 and Type 2 parameter sets, identified as the MCS. Our calculation results suggest that these plateaus could potentially be verified experimentally through magnetization measurements under the magnetic field of approximately 300 T, achievable with state-of-the-art magnetic field generators [38, 39, 40]. Notably, clear 5/9 and 7/9 magnetization plateaus have not yet been observed in experiments, highlighting the importance of additional experimental investigation into $\rm{Y}_{3}\rm{Cu}_{9}(\rm{OH})_{19}\rm{Cl}_{8}$. Additionally, by performing calculations of the kagome lattice with parameter sets other than the Type 1 and Type 2, we discovered additional plateau phases at $M/M_{\rm{sat}}=1/3$, 5/9, and 7/9. The mechanism behind the 1/3 plateau can be easily explained by analytically solving a three-site system. In contrast, the mechanisms for the 5/9 and 7/9 plateaus could not be clarified in this study. Further experimental and theoretical research is therefore needed to elucidate these phenomena. These findings indicate that further experimental studies on new kagome compounds may successfully observe these new plateaus.